1. A. Determine the number of atoms in 3.50 mol Fe (iron). B. Calculate the number of molecules in 12.5 mol H₂O 2. Determine the number of moles in each of the ff. A. 125g Zn (zinc) B.1.00 kg S (sulfur) 3. How many chloride ions (Cl_) are present in 2.50 moles of CaCl2( calcium chloride) 4.convert 5.12 mol Of sulfuric acid H₂SO₄
1
Expert's answer
2021-10-28T03:00:51-0400

Molar mass of Fe = 55.845

3.5/55.845= 0.062atoms

1 mole of H2O = 6.022× 1023molecules

6.022×1023×12.5= 7.53×1024molecules